Technical Staff Blog

Category archives: Service Outages

Tstaff announcement about complete service outages

RSS feed of Service Outages

Last update on .

The department list server stopped delivering mail starting at least as far back as yesterday morning (Thursday the 10th). The problem has been resolved and the list server is chewing through the backlog of messages. If you continue to have problems sending email to department lists or receiving listserv messages, email problem.

Last update on .

CIS will be performing a planned switch upgrade on the 3rd floor of the CIT on Thursday, 1/10 from 6:00 AM to 8:00 AM. During this time users on the 3rd floor may experience brief interruptions (less than 5 minutes) of wired, wireless network connectivity and potentially phone usage while the connections ...

Last update on .

CIS will be performing a planned switch upgrade on the 5th floor of the CIT on Tuesday, 1/8 from 6:00 AM to 9:00 AM. During this time users on the 4th and 5th floors may experience brief interruptions (less than 5 minutes) of wired, wireless network connectivity and potentially phone usage while ...

Last update on .

Many services were affected by a GPFS filesystem issue this morning.  An NFS server in our GPFS cluster failed, affecting filesystem access generally, and services that rely on filesystem access.  Systems were affected by at least 10am, and possibly much earlier.  The filesystem was restored to normal operation before 2pm.  A VMWare host server also ...

Last update on .

Reports came in yesterday and today of messages being delayed for many hours.  Some were specifically because of account name changes, which was resolved early in the day, yesterday.  While others were delayed due to an after hours service outage, which has also been resolved.

Email services have been returned to normal and all delayed ...

Last update on .

This is an after the fact report - this issue has been resolved. Certain parts of the 4th and 5th floor of CIT experienced a brief network outage on Thursday, October 4th sometime between 2pm and 3pm. TStaff immediately contacted CIS Networking and they dispatched Comm Ops, it appeared to be a bad uplink on one ...

Last update on .

Users have reported an unusual volume and variety of FastX problems over the last couple of days. So far we have been unable to determine the precise cause, or narrow the problems down to any particular host in the cluster. Therefore, as part of our ongoing attempt to diagnose and fix the problems, we are ...

Last update on .

A number of services were intermittently unavailable today.  These include VPN, the list server and the website.  The cause of the problems is not yet known and we continue to investigate.  There remain lingering problems - intermittent authentication failures on the website in particular - and we have posted a notice on the system status page.

Last update on .

All critical CS Department services have been restored. An interruption in external services resulted in a cascading DNS failure and overall service instability, starting mid-morning on Thursday of this week.  After repeated efforts to diagnose and workaround the issue, we reached out to CIS who stuck with us throughout today to diagnose. Ultimately, we were ...